# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and variable-order nonlocal operators

### v2

@@ 2.4. Solvability of semilinear stochastic integro-diffrential equations
where $\mathcal{D}^{L}_{h}u(x):=\mathcal{D}_{h}(\mathcal{D}^{L-1}_{h}u)(x)$, $\mathcal{D}_{h}u(x):=u(x+h)-u(x)$, $L>\gamma-d/p$, and $\delta_{0}\gamma-d/p=n+\nu$ for $\nu\in(0,1]$. By applying (3.24) and considering the continuous embedding given by Theorem 3.7, i.e.,
@@
$\geq N\mathbb{E}\left[\left(\sup_{|h|\leq 1}\frac{|\mathcal{D}_{h}^{L}u|_{C([0,T]\times\mathbb{R}^{d})}}{\phi(|h|^{-2})^{-\gamma+2\beta}|h|^{-d/p}}+\sup_{0\leq s<t\leq T}\frac{|u(t,\cdot)-u(s,\cdot)|_{C(\mathbb{R}^{d})}}{(t-s)^{\alpha-1/q}}\right)^{q}\right]$
@@
One can check that
$\frac{|\mathcal{D}_{(h_{1},h_{2})}^{L}u|_{C([0,T]\times\mathbb{R}^{d})}}{|h_{1}|^{\alpha-\frac{1}{q}}+\phi(|h_{2}|^{-2})^{-\gamma+2\beta}|h_{2}|^{-\frac{d}{p}}}$
$\leq N\sum_{i=1}^{L}\frac{|u(\cdot+ih_{1},\cdot)-u(\cdot+(i-1)h_{1},\cdot)|_{C([0,T]\times\mathbb{R}^{d})}}{|h_{1}|^{\alpha-\frac{1}{q}}}+N\frac{|\mathcal{D}_{h_{2}}^{L}u|_{C([0,T]\times\mathbb{R}^{d})}}{\phi(|h_{2}|^{-2})^{-\gamma+2\beta}|h_{2}|^{-\frac{d}{p}}}.$
These estimations lead us to conclude that
@@
22 22 $\geq\frac{1}{A}\int_{0}^{d_{n-1}}\int_{0}^{1}\frac{\|h(t)-h(s)\|_{V}^{q}}{|t-s|^{\alpha q+1}}\mathrm{d}t\mathrm{d}s+\frac{1}{B}\int_{0}^{d_{n-1}}\frac{\|h(t_{n-1})-h(s)\|_{V}^{q}}{|t_{n-1}-s|^{\alpha q+1}}\mathrm{d}s$
@@
Initially, we utilize the isometry to set $\gamma=2\beta$. Given $\tau\leq T$ and $u\in\mathcal{H}_{p,q}^{\phi,\gamma}(\tau)$, we define $f:=(\mathbb{D}u-\phi(\Delta)u)1_{t\leq\tau}$ and $g=\mathbb{S}u1_{t\leq\tau}$. Consequently, $f$ resides within $\mathbb{H}_{p,q}^{\phi,\gamma-2}(T)$, and $g$ within $\mathbb{H}_{p,q}^{\phi,\gamma-1}(T,\ell_{2})$. Leveraging Theorem 2.10, we establish that:
@@
on  |($0,\tau_{m}(S)\text{$]$\kern-1.49994pt$]$}\times\mathbb{R}^{d}$ with initial data $u_{0}$.
@@
$\frac{1}{p}<\frac{d+1-\delta_{0}(2-\gamma)}{d}<\frac{1}{r}<1+\frac{1}{p}+\frac{\lambda_{\mathrm{s.d.}}}{p\vee q}-\lambda_{\mathrm{b.}}.$
@@
$10<\frac{1}{r}+\lambda_{\mathrm{b.}}-\frac{1}{p}-\frac{\lambda_{\mathrm{s.d.}}}{p\vee q}<$ 1 . (3.42)
Using (1.5), we also have
$\|R_{\delta_{0}(2-\gamma)-1}\|_{L_{r}(\mathbb{R}^{d})}<\infty.$
If $q\leq p$, then by Hölder’s inequality,
@@
×$\qquad\times\left(\int_{0}^{\tau_{m}(S)}\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{1+\lambda_{\mathrm{s.d.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\mathrm{d}t\right)^{\frac{q}{p}}$
@@
×$\qquad\times\left(\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{\frac{p}{r}+p\lambda_{\mathrm{b.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\right)\mathrm{d}t$
@@
×$\qquad\times\int_{0}^{\tau_{m}(S)}\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{1+\lambda_{\mathrm{s.d.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\mathrm{d}t$
@@
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}R_{\delta_{0}(1-\gamma)}(x-(y-z))R_{\delta_{0}(1-\gamma)}(x-z)f(y-z)\overline{f(-z)}\mathrm{d}z\mu(\mathrm{d}y).$
@@
$\leq\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|u_{m}(t,y-z)|^{\frac{1}{2}}|u_{m}(t,-z)|^{\frac{1}{2}}\mathrm{d}z\frac{\mu(\mathrm{d}y)}{(1+|y|^{2})^{k/2}}\right)^{1-\frac{1}{r}}$
×$\quad\times\bigg{(}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|R_{\delta_{0}(1-\gamma)}(x-(y-z))R_{\delta_{0}(1-\gamma)}(x-z)|^{r}$
×$\quad\quad\quad\times|u_{m}(t,y-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}|u_{m}(t,-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}$
×$\quad\quad\quad\times h_{m}(u_{m}(t,y-z))h_{m}(u_{m}(t,-z))(1+|y|^{2})^{k(r-1)/2}\mathrm{d}z\mu(\mathrm{d}y)\bigg{)}^{\frac{1}{r}}$
$\leq A_{\mu}^{1-\frac{1}{r}}\|u_{m}(t,\cdot)\|_{L_{1}(\mathbb{R}^{d})}^{1-\frac{1}{r}}\Gamma(x)^{\frac{1}{r}}\leq A_{\mu}^{1-\frac{1}{r}}S^{1-\frac{1}{r}}\Gamma(x)^{\frac{1}{r}},$
where $A_{\mu}$ is the constant defined in (1.10) and
@@
×$\quad\quad\quad\times|u_{m}(t,y-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}|u_{m}(t,-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}$
×$\quad\quad\quad\times h_{m}(u_{m}(t,y-z))h_{m}(u_{m}(t,-z))(1+|y|^{2})^{k(r-1)/2}\mathrm{d}z\mu(\mathrm{d}y).$
@@
$\int_{\mathbb{R}^{d}}\Gamma(x)^{\frac{p}{2r}}\mathrm{d}x$ $\leq\left(\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(y)(1+|y|^{2})^{k(r-1)/2}\mu(\mathrm{d}y)\right)^{\frac{p}{2r}}$
×$\quad\times\left(\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{\frac{p(r+1)}{2r}+p\lambda_{\mathrm{s.m.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\right)$
@@
$I<\infty\Longleftrightarrow\int_{|y|<1}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(y)\mu(\mathrm{d}y)<\infty.$
@@
By making use of Chebyshev’s inequality, Corollary 3.6 and (3.43),
@@
Acknowledgements. J.-H. Choi has been supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(ME) (No.RS-2023-00237315). B.-S. Han has been supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2021R1C1C2007792). D. Park has been supported by the 2024 Research Grant from Kangwon National University (grant no. 202404630001).

### v3

@@ 2.4. Solvability of semilinear stochastic integro-differential equations
where $\mathcal{D}^{L}_{h}u(x):=\mathcal{D}_{h}(\mathcal{D}^{L-1}_{h}u)(x)$, $\mathcal{D}_{h}u(x):=u(x+h)-u(x)$, $L$ is an integer greater than $\gamma-d/p$, and $\delta_{0}\gamma-d/p=n+\nu$ for $n\in\mathbb{N}\cup\{0\}$ and $\nu\in(0,1]$. By applying (3.31) and considering the continuous embedding given by Theorem 3.12, i.e.,
@@
$\geq N\mathbb{E}\left[\left(\sup_{|h|\leq 1}\frac{|\mathcal{D}_{h}^{L}u|_{C([0,T]\times\mathbb{R}^{d})}}{\phi(|h|^{-2})^{-\frac{\gamma}{2}+\beta}|h|^{-d/p}}+\sup_{0\leq s<t\leq T}\frac{|u(t,\cdot)-u(s,\cdot)|_{C(\mathbb{R}^{d})}}{(t-s)^{\alpha-1/q}}\right)^{q}\right]$
@@
Proposition 3.9 leads us to conclude that
@@
$2$ $\geq\frac{1}{A}\int_{0}^{d_{n-1}}\int_{0}^{1}\frac{\|h(t)-h(s)\|_{V}^{q}}{|t-s|^{\alpha q+1}}\mathrm{d}t\mathrm{d}s+\frac{1}{B}\int_{0}^{d_{n-1}}\frac{\|h(t_{n-1})-h(s)\|_{V}^{q}}{|t_{n-1}-s|^{\alpha q+1}}\mathrm{d}s$
@@
It suffices to prove the desired result for the case $\gamma=2\beta$, since the general case $\gamma\in\mathbb{R}$ can be reduced to this special case via the isometry induced by the operator $(1-\phi(\Delta))^{\frac{\gamma}{2}-\beta}$. Indeed, we have
$N\|u\|_{\mathcal{H}_{p,q}^{\phi,\gamma}(\tau)}^{q}$
$=N\|(1-\phi(\Delta))^{\frac{\gamma}{2}-\beta}u\|_{\mathcal{H}_{p,q}^{\phi,2\beta}(\tau)}^{q}$
$\geq\mathbb{E}\left[\sup_{t\leq\tau}\|(1-\phi(\Delta))^{\frac{\gamma}{2}-\beta}u(t,\cdot)\|_{L_{p}(\mathbb{R}^{d})}^{q}\right]$
$\quad+\mathbb{E}\left[\left(\sup_{0\leq s<t\leq\tau}\frac{\|((1-\phi(\Delta))^{\frac{\gamma}{2}-\beta}(u(t,\cdot)-u(s,\cdot))\|_{L_{p}(\mathbb{R}^{d})}}{(t-s)^{\alpha-1/q}}\right)^{q}\right]$
$=\mathbb{E}\left[\sup_{t\leq\tau}\|u(t,\cdot)\|_{H_{p}^{\gamma-2\beta}(\mathbb{R}^{d})}^{q}\right]+\mathbb{E}\left[\left(\sup_{0\leq s<t\leq\tau}\frac{\|u(t,\cdot)-u(s,\cdot)\|_{H_{p}^{\phi,\gamma-2\beta}(\mathbb{R}^{d})}}{(t-s)^{\alpha-1/q}}\right)^{q}\right].$
Therefore, it is enough to establish the estimate for $\gamma=2\beta$.
Given $\tau\leq T$ and $u\in\mathcal{H}_{p,q}^{\phi,\gamma}(\tau)$, we define $f:=(\mathbb{D}u-\phi(\Delta)u)1_{t\leq\tau}$ and $g=\mathbb{S}u1_{t\leq\tau}$. Consequently, $f$ resides within $\mathbb{H}_{p,q}^{\phi,\gamma-2}(T)$, and $g$ within $\mathbb{H}_{p,q}^{\phi,\gamma-1}(T,\ell_{2})$. Leveraging Theorem 2.10, we establish that:
@@
[…]
@@
$\frac{1}{p}<\frac{d+1-\delta_{0}(2-\gamma)}{d}<\frac{1}{r}<1+\frac{1}{p}+\frac{\lambda_{\mathrm{s.d.}}}{p\vee q}-\lambda_{\mathrm{b.}}.$ (3.49)
@@
$0<\frac{1}{r}+\lambda_{\mathrm{b.}}-\frac{1}{p}-\frac{\lambda_{\mathrm{s.d.}}}{p\vee q}<1.$ (3.50)
Since $\delta_{0}(2-\gamma)-1\in(0,1)$, it follows from (1.5) and (1.6) that
$\|R_{\delta_{0}(2-\gamma)-1}\|_{L_{r}(\mathbb{R}^{d})}<\infty\quad\Longleftrightarrow\quad r<\frac{d}{d+1-\delta_{0}(2-\gamma)}.$
This condition is clearly satisfied by the choice of $r$ in (3.49). If $q\leq p$, then by Hölder’s inequality,
@@
$\qquad\times\left(\int_{0}^{\tau_{m}(S)}\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{1+\lambda_{\mathrm{s.d.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\mathrm{d}t\right)^{\frac{q}{p}}$
@@
$\qquad\times\left(\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{\frac{p}{r}+p\lambda_{\mathrm{b.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\right)\mathrm{d}t$
@@
$\qquad\times\int_{0}^{\tau_{m}(S)}\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{1+\lambda_{\mathrm{s.d.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\mathrm{d}t$
@@
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}R_{\delta_{0}(1-\gamma)}(x-(y-z))R_{\delta_{0}(1-\gamma)}(x+z)f(y-z)\overline{f(-z)}\mathrm{d}z\pi(\mathrm{d}y).$
@@
$\leq\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|u_{m}(t,y-z)|^{\frac{1}{2}}|u_{m}(t,-z)|^{\frac{1}{2}}\mathrm{d}z\frac{\pi(\mathrm{d}y)}{(1+|y|^{2})^{k/2}}\right)^{1-\frac{1}{r}}$
$\quad\times\bigg{(}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|R_{\delta_{0}(1-\gamma)}(x-(y-z))R_{\delta_{0}(1-\gamma)}(x-z)|^{r}$
$\quad\quad\quad\times|u_{m}(t,y-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}|u_{m}(t,-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}$
$\quad\quad\quad\times h_{m}(u_{m}(t,y-z))h_{m}(u_{m}(t,-z))(1+|y|^{2})^{k(r-1)/2}\mathrm{d}z\pi(\mathrm{d}y)\bigg{)}^{\frac{1}{r}}$
$\leq A_{\pi}^{1-\frac{1}{r}}\|u_{m}(t,\cdot)\|_{L_{1}(\mathbb{R}^{d})}^{1-\frac{1}{r}}\Gamma(x)^{\frac{1}{r}}\leq A_{\pi}^{1-\frac{1}{r}}S^{1-\frac{1}{r}}\Gamma(x)^{\frac{1}{r}},$
where $A_{\pi}$ is the constant defined in (1.11) and
@@
$\quad\quad\quad\times|u_{m}(t,y-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}|u_{m}(t,-z)|^{\frac{r+1}{2}+r\lambda_{\mathrm{s.m.}}}$
$\quad\quad\quad\times h_{m}(u_{m}(t,y-z))h_{m}(u_{m}(t,-z))(1+|y|^{2})^{k(r-1)/2}\mathrm{d}z\pi(\mathrm{d}y).$
@@
$\int_{\mathbb{R}^{d}}\Gamma(x)^{\frac{p}{2r}}\mathrm{d}x$ $\leq\left(\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(y)(1+|y|^{2})^{k(r-1)/2}\pi(\mathrm{d}y)\right)^{\frac{p}{2r}}$
$\quad\times\left(\int_{\mathbb{R}^{d}}|u_{m}(t,x)|^{\frac{p(r+1)}{2r}+p\lambda_{\mathrm{s.m.}}}h_{m}(u_{m}(t,x))\mathrm{d}x\right)$
@@
$I<\infty\Longleftrightarrow\int_{|y|<1}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(y)\pi(\mathrm{d}y)<\infty.$
@@
By making use of Chebyshev’s inequality, Theorem 3.11 and (3.51),
@@
Acknowledgements. The authors are highly indebted to the anonymous reviewers for their careful reading and insightful comments, which substantially improved the presentation of this paper. J.-H. Choi has been supported by a KIAS Individual Grant (MG102701) at Korea Institute for Advanced Study. B.-S. Han has been supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2021R1C1C2007792). D. Park has been supported by the 2025 Research Grant from Kangwon National University (grant no. 202504300001). The authors would also like to thank Dr. Junhee Ryu and Dr. Jinsol Seo for giving many helpful comments.